West Coast Utilities had a net profit of $933 million. It has 911million shares outstanding and paid annual dividends of$ 0.58 per share. What is the dividend payout ratio?

The dividend payout ratio is % (Round to two decimal places.)
Answer #1

dividend payout ratio = dividend per share / earning per share

dividend per share = 0.58

earning per share = net profit / shares outstanding = 933 m / 911 m = 1.02415

dividend payout ratio = 0.58 / 1.02415 = 56.63%

Hence the answer is 56.63%